### v2

@@ 1 Introduction
Recent literature has seen a burgeoning interest in the study of extreme outcomes, exemplified by frameworks such as Financial Stability at Risk as measured by CoVaR (adrian2016covar), Employment at Risk (ramey2018government), Growth at Risk (adrian2019vulnerable), Inflation at Risk (loria2022inflation), and the Term Structure of Growth at Risk (adrian2022term), to name but a few. Given the scarcity of data in the tails of outcome distributions, conventional econometric methods are often inadequate for these analyses. To address this challenge, researchers have developed novel econometric techniques tailored specifically to evaluate and quantify risks associated with such extreme outcomes.
To fix ideas, we focus on the Growth-at-Risk (GaR) framework of adrian2019vulnerable. Let $Y_{t+h}$ denote the average annual GDP growth rate between periods $t$ and $t+h$. Let $X_{t}$ denote a vector of observed controls. Researchers are interested in measuring the tail features of the conditional distribution of $Y_{t+h}$ given $X_{t}$. For instance, a primary object of interest is the conditional quantile
@@
While the existing literature offers several valuable methods for analyzing these measures, it also has notable limitations. For example, if one begins with linear quantile regressions $Q_{Y_{t+h}|X_{t}}(\tau|x)=x^{\prime}\beta(\tau)$ and extrapolates them to the tails using conditional skewed-t or Pareto-Lévy distributions (mandelbrot1960pareto) of $Y_{t+h}$ given $X_{t}=x$, then the resulting estimation strategy implicitly assumes a conditional Pareto exponent (equivalent to the degree of freedom of Student-t) that remains constant with respect to $x$. Consequently, the expected shortfall $SF_{Y_{t+h}|X_{t}}(\pi|x)$ and expected longrise $LR_{Y_{t+h}|X_{t}}(\pi|x)$ become proportional to the baseline linear quantile regression $Q_{Y_{t+h}|X_{t}}(\tau|x)=x^{\prime}\beta(\tau)$. In this paper, we will theoretically demonstrate these results.
Given the limitations of existing methods, we propose a more robust econometric approach for analyzing these parameters. Unlike existing methods, our approach does not imply a constant Pareto exponent with respect to $x$. We provide a formal asymptotic theoretical foundation to validate our proposed method. Additionally, simulation studies based on data-generating processes that mimic the real-world data used in adrian2019vulnerable demonstrate that our method consistently outperforms existing alternatives in terms of prediction accuracy.
@@ 2.1 Review of the Existing Method
We begin with reviewing a popular existing econometric methods. While we highlight the work of adrian2019vulnerable due to its prominence, with other existing approaches exhibiting similar underlying characteristics. adrian2019vulnerable propose the following estimation strategy.
@@ 2.2.1 Constancy of $v(x)$ in $x$
This result originates from wang2013estimation, who study the constancy of the tail index. We tailor their argument for the skewed-t distribution in our setup. A proof is found in Appendix A.1. This proposition establishes that the linear quantile regression model $Q_{Y_{t+h}|X_{t}}(x)=x^{\prime}\beta(\tau)$ assumed in Step 1 of Section 2.1 and the skewed-t distribution of $F_{Y_{t+h}|X_{t}}\left(\ \cdot\ |x\right)$ assumed in Step 2 of Section 2.1 necessarily imply that degree of freedom $v(x)=\overline{v}$ is constant and does not depend on observed predictor $x$. In summary, while the procedure of adrian2019vulnerable gives a tail index $v(x)$ that depends on $x$, their model implies that the tail in fact does not depend on $x$ but is a constant $v(x)=\overline{v}$. Proposition 1 implies an internal inconsistency within the existing method: fitting the skewed-t distribution and using quantile regression are self-contradicting. Specifically, quantile regression implies that the conditional quantile depends on $x$ through a linear location-and-scale transformation. Such location-and-scale model restricts that the tail index, which corresponds to the degree of freedom in the skewed-t distribution, to be constant. Therefore, fitting the skewed-t distribution automatically loses such dependence when using extreme quantiles. Since GaR precisely targets the extreme quantile and other tail features, losing the dependence of tail shape on $x$ could incur substantial misspecification bias.
@@ 3 Our New Econometric Method of Growth at Risk
for a sufficiently large $y_{\min}$, where $v(x)$ denotes the tail exponent. We emphasize that the skewed-t distribution employed by adrian2019vulnerable satisfies this approximation condition (1) with $v(x)$ corresponding to the degrees of freedom. Besides the skewed-t distribution, it accommodates many parametric and nonparametric families (e.g., dehaan2006extreme, Chapter 1). Hence, this assumption (1) is fairly weak and much weaker than the implicit assumptions made in the prior literature.
@@
for example. The log-linear form has been explored by the statistics literature (e.g., wang2009tail; wang2013estimation) with two major benefits: (i) it guarantees a positive tail index, and (ii) the tail index regression problem in Step 1 is strictly convex. We thus consider this functional form as a parsimonious and convenient specification. Again, we want to emphasize that this model (2) is much more general than the implicit assumption imposed in the prior literature that $v(x)=\overline{v}$ is constant in $x$.
@@ 3.1.1 Algorithm for the Upper Tail
We provide some discussions. Step 1 is the tail index regression proposed by wang2009tail. We modify their approach by subtracting the sample median $\overline{Y}$, which will not change the asymptotic distribution. The tail cutoff $y_{\min}$ serves as a tuning parameter, close in spirit to the bandwidth in kernel regression. We use 90% quantile as a simple rule-of-thumb choice here and present a more sophisticated data-driven choice in Section 3.3. Once we obtain the tail index estimator $\hat{v}(x_{0})$, we now take advantage of the Pareto tail approximation to estimate in the conditional extreme quantile as in Step 3. The expected longrise can be estimated analogously.
@@ 3.3 Rule-of-Thumb Choice of the Threshold
The choice of $y_{\min}$ (and $y_{\max}$) is similar to that of the bandwidth in kernel regression. There is a large literature in studying the optimal choice of $y_{\min}$ in the case without the covariate. See, danielsson2001using; danielsson2016tail, among many other. In the case with covariate, we provide a data-driven method to choose the threshold, following wang2009tail. For simple illustration, we focus on the right tail.
@@ 4 Asymptotic Theory
Assumption P imposes a second-order conditional Pareto tail, with the term $y^{-\rho(x_{0})}$ characterizing the deviation from the exact Pareto – see also wang2009tail. Assumption N requires that the nonparametric estimator $\hat{F}_{Y_{t+h}|X_{t}}\left(\cdot|x_{0}\right)$ uniformly converges to a Gaussian process with the variance-covariance proportional to $1-F_{Y_{t+h}|X_{t}}(\cdot|x_{0})$ in the right tail. Specifically, this can be established for kernel estimation with $r_{T}=b^{\text{dim}(X)/2}_{T}$, where $b_{T}$ is the bandwidth. The constant $c(x_{0})$ then becomes $\int K(u)^{2}du/g(x_{0})$ where $g(\cdot)$ denotes the density of $X$ and $K(\cdot)$ the kernel function (e.g., cai2002regression; li2008nonparametric). [^3] These papers establish the pointwise convergence of the kernel conditional CDF estimator. Extension to the uniform convergence follows from the standard empirical process argument (e.g., arcones1994central). Such a result is available upon request. Assumption T describes the convergence of the tail index regression estimator. This condition is established by nicolau2023tail and wang2009tail for the case without subtracting $\bar{Y}$. Suppose $\bar{Y}$ converges to the median of $Y$ at the rate root-$T$, subtracting $\bar{Y}$ from the extreme values of $Y$ makes no asymptotic effect.
@@
Suppose Assumptions P, N, T, R hold. Then,
@@ 5.1 Data Generating Designs
We develop two data generating designs based on the real data used by adrian2019vulnerable. These two designs are tailored to quarter-ahead predictions (Section 5.1.1) and year-ahead predictions (Section 5.1.2) of GDP growth rates, based on the current GDP growth rate and a financial indicator.
@@ 5.1.1 Design Based on Quarter-Ahead Predictions
These parameter values are obtained by fitting the known model of the conditional skewed-t distribution with the real data used by adrian2019vulnerable for quarter-ahead predictions.
@@ 5.1.2 Design Based on Year-Ahead Predictions
These parameter values are obtained by fitting the known model of the conditional skewed-t distribution with the real data used by adrian2019vulnerable for year-ahead predictions.
@@ 5.4 Additional Simulations
Having demonstrated the superior performance of our proposed method in predicting extreme growth events, we now apply this novel approach to a long-term U.S. macro-financial dataset spanning 1893–2016. The dataset is obtained from Gachter_Hasler_Huber_2025 stretching from 1893Q1 to 2016Q4 for the US and including annualized real GDP growth, a financial stress indicator, the 3-year average growth rate of the credit-to-GDP ratio, and the 3-year average growth rate of real house prices. The financial stress indicators is a historical newspaper-based financial stress indicator.
@@ 7 Summary and Discussions
While our primary focus is on the GaR framework, the proposed method is broadly applicable to other contexts involving the study of extreme outcomes, including Financial Stability at Risk (adrian2016covar), Employment at Risk (ramey2018government), Growth at Risk (adrian2019vulnerable), Inflation at Risk (loria2022inflation), and Term Structure of Growth at Risk (adrian2022term), among others. We hope this paper will inspire further empirical research on these and other topics in economics and finance.
@@ Appendix A Mathematical Proofs
This DoA condition holds for a broad class of nonparametric distributions, including the skewed-t in particular. Thus, adrian2019vulnerable implicitly assume this DoA condition too.
@@
This RV condition holds for a broad class of nonparametric distributions, including the skewed-t in particular. Thus, adrian2019vulnerable implicitly assume this RV condition too.
@@ A.1 Proof of Proposition 1
Proof. For simplicity of writing, we abbreviate the notations $Y_{t+h}$ and $X_{t}$ as $Y$ and $X$, respectively, in this proof. We focus on the right tail, as the argument for the left tail follows symmetrically. Since the skewed-t distribution satisfies the DoA assumption with $\xi\left(x\right)=v\left(x\right)>0$, it follows that $Q_{Y|X}\left(\tau|x\right)\rightarrow\infty$ as $\tau\rightarrow 1$ and $Q_{Y|X}\left(\cdot|x\right)$ is regularly varying with coefficient $v\left(x\right)>0$.
Let $x=(1,\dot{x}^{\prime})^{\prime}$ so that $Q_{Y|X}\left(\tau|x\right)=\alpha(\tau)+\dot{x}^{\prime}\beta(\tau)$. Let $j^{*}$ denote the coordinate of $\beta$ that has the heaviest tail, i.e., $\lim_{\tau\to 1}\beta_{j}(\tau)/\beta_{j^{*}}(\tau)=0$ for all $j\neq j^{*}$. [^4] If two components have the same order, i.e., $\lim_{\tau\to 1}\beta_{j}(\tau)/\beta_{j^{*}}(\tau)=c$ for some constant $c\in(0,\infty)$ and some $j$, then combine components $j$ and $j^{*}$ and consider $x_{j^{*}}+cx_{j}$ to be the dominating component. Now let $h=\lim_{\tau\to 1}\alpha(\tau)/\beta(\tau)>0$. We now proceed with each of the three cases, $h=0$, $h=\infty$, and $h\in(0,\infty)$, separately.
Case 1: $h=0$. In this case, $\beta(1-1/t)=o(\alpha(1-1/t))$ as $t\rightarrow\infty$ and hence
@@
Case 3: $h\in(0,\infty)$. In this case, $\alpha(1-1/t)\sim h\beta_{j^{*}}(1-1/t)$ as $t\rightarrow\infty$ and hence
$\lim_{t\to\infty}\frac{Q_{Y|X}\left(1-\frac{1}{t\tau}|x\right)}{Q_{Y|X}\left(1-\frac{1}{t}|x\right)}=\lim_{t\to\infty}\frac{(\dot{x}_{j^{*}}+h)\beta_{j^{*}}\left(1-\frac{1}{t\tau}\right)}{(\dot{x}_{j^{*}}+h)\beta_{j^{*}}\left(1-\frac{1}{t}\right)}=\lim_{t\to\infty}\frac{\beta_{j^{*}}\left(1-\frac{1}{t\tau}\right)}{\beta_{j^{*}}\left(1-\frac{1}{t}\right)},$
@@
Combining three cases yields that $\lim_{t\to\infty}\frac{Q_{Y|X}\left(1-\frac{1}{t\tau}|x\right)}{Q_{Y|X}\left(1-\frac{1}{t}|x\right)}$ does not depend on $\dot{x}$. However, by regular varying, we have
$\frac{Q_{Y|X}\left(1-\frac{1}{t\tau}|x\right)}{Q_{Y|X}\left(1-\frac{1}{t}|x\right)}\rightarrow\tau^{v\left(x\right)}$

### v3

@@ 1 Introduction
Recent literature has seen a burgeoning interest in the study of extreme outcomes, exemplified by frameworks such as Financial Stability at Risk as measured by CoVaR (Adrian and Brunnermeier, 2016), Employment at Risk (Ramey and Zubairy, 2018), Growth at Risk (Adrian et al., 2019), Inflation at Risk (Loria et al., 2022), and the Term Structure of Growth at Risk (Adrian et al., 2022), to name but a few. Given the scarcity of data in the tails of outcome distributions, conventional econometric methods are often inadequate for these analyses. To address this challenge, researchers have developed novel econometric techniques tailored specifically to evaluate and quantify risks associated with such extreme outcomes.
To fix ideas, we focus on the Growth-at-Risk (GaR) framework of Adrian et al. (2019). Let $Y_{t+h}$ denote the average annual GDP growth rate between periods $t$ and $t+h$. Let $X_{t}$ denote a vector of observed controls. Researchers are interested in measuring the tail features of the conditional distribution of $Y_{t+h}$ given $X_{t}$. For instance, a primary object of interest is the conditional quantile
@@
While the existing literature offers several valuable methods for analyzing these measures, it also has notable limitations. For example, if one begins with linear quantile regressions $Q_{Y_{t+h}|X_{t}}(\tau|x)=x^{\prime}\beta(\tau)$ and extrapolates them to the tails using conditional skewed-t or Pareto-Lévy distributions (Mandelbrot, 1960) of $Y_{t+h}$ given $X_{t}=x$, then the resulting estimation strategy implicitly assumes a conditional Pareto exponent (equivalent to the degree of freedom of Student-t) that remains constant with respect to $x$. Consequently, the expected shortfall $SF_{Y_{t+h}|X_{t}}(\pi|x)$ and expected longrise $LR_{Y_{t+h}|X_{t}}(\pi|x)$ become proportional to the baseline linear quantile regression $Q_{Y_{t+h}|X_{t}}(\tau|x)=x^{\prime}\beta(\tau)$. In this paper, we will theoretically demonstrate these results.
Given the limitations of existing methods, we propose a more robust econometric approach for analyzing these parameters. Unlike existing methods, our approach does not imply a constant Pareto exponent with respect to $x$. We provide a formal asymptotic theoretical foundation to validate our proposed method. Additionally, simulation studies based on data-generating processes that mimic the real-world data used in Adrian et al. (2019) demonstrate that our method consistently outperforms existing alternatives in terms of prediction accuracy.
@@ 2.1 Review of the Existing Method
We begin with reviewing a popular existing econometric methods. While we highlight the work of Adrian et al. (2019) due to its prominence, with other existing approaches exhibiting similar underlying characteristics. Adrian et al. (2019) propose the following estimation strategy.
@@ 2.2.1 Constancy of $v(x)$ in $x$
This result originates from Wang and Li (2013, Proposition 2.1), who study the constancy of the tail index. We tailor their argument for the skewed-t distribution in our setup. A proof is found in Appendix A.1. This proposition establishes that the linear quantile regression model $Q_{Y_{t+h}|X_{t}}(x)=x^{\prime}\beta(\tau)$ assumed in Step 1 of Section 2.1 and the skewed-t distribution of $F_{Y_{t+h}|X_{t}}\left(\ \cdot\ |x\right)$ assumed in Step 2 of Section 2.1 necessarily imply that degree of freedom $v(x)=\overline{v}$ is constant and does not depend on observed predictor $x$. In summary, while the procedure of Adrian et al. (2019) gives a tail index $v(x)$ that depends on $x$, their model implies that the tail in fact does not depend on $x$ but is a constant $v(x)=\overline{v}$. Proposition 1 implies an internal inconsistency within the existing method: fitting the skewed-t distribution and using quantile regression are self-contradicting. Specifically, quantile regression implies that the conditional quantile depends on $x$ through a linear location-and-scale transformation. Such location-and-scale model restricts that the tail index, which corresponds to the degree of freedom in the skewed-t distribution, to be constant. Therefore, fitting the skewed-t distribution automatically loses such dependence when using extreme quantiles. Since GaR precisely targets the extreme quantile and other tail features, losing the dependence of tail shape on $x$ could incur substantial misspecification bias.
@@ 3 Our New Econometric Method of Growth at Risk
for a sufficiently large $y_{\min}$, where $v(x)$ denotes the tail exponent. We emphasize that the skewed-t distribution employed by Adrian et al. (2019) satisfies this approximation condition (1) with $v(x)$ corresponding to the degrees of freedom. Besides the skewed-t distribution, it accommodates many parametric and nonparametric families (e.g., de Haan and Ferreira, 2006, Chapter 1). Hence, this assumption (1) is fairly weak and much weaker than the implicit assumptions made in the prior literature.
@@
for example. The log-linear form has been explored by the statistics literature (e.g., Wang and Tsai, 2009; Wang and Li, 2013) with two major benefits: (i) it guarantees a positive tail index, and (ii) the tail index regression problem in Step 1 is strictly convex. We thus consider this functional form as a parsimonious and convenient specification. Again, we want to emphasize that this model (2) is much more general than the implicit assumption imposed in the prior literature that $v(x)=\overline{v}$ is constant in $x$.
@@ 3.1.1 Algorithm for the Upper Tail
We provide some discussions. Step 1 is the tail index regression proposed by Wang and Tsai (2009). We modify their approach by subtracting the sample median $\overline{Y}$, which will not change the asymptotic distribution. The tail cutoff $y_{\min}$ serves as a tuning parameter, close in spirit to the bandwidth in kernel regression. We use 90% quantile as a simple rule-of-thumb choice here and present a more sophisticated data-driven choice in Section 3.3. Once we obtain the tail index estimator $\hat{v}(x_{0})$, we now take advantage of the Pareto tail approximation to estimate in the conditional extreme quantile as in Step 3. The expected longrise can be estimated analogously.
@@ 3.3 Rule-of-Thumb Choice of the Threshold
The choice of $y_{\min}$ (and $y_{\max}$) is similar to that of the bandwidth in kernel regression. There is a large literature in studying the optimal choice of $y_{\min}$ in the case without the covariate. See, Danielsson et al. (2001, 2016), among many other. In the case with covariate, we provide a data-driven method to choose the threshold, following Wang and Tsai (2009). For simple illustration, we focus on the right tail.
@@ 4 Asymptotic Theory
Assumption P imposes a second-order conditional Pareto tail, with the term $y^{-\rho(x_{0})}$ characterizing the deviation from the exact Pareto – see also Wang and Tsai (2009). Assumption N requires that the nonparametric estimator $\hat{F}_{Y_{t+h}|X_{t}}\left(\cdot|x_{0}\right)$ uniformly converges to a Gaussian process with the variance-covariance proportional to $1-F_{Y_{t+h}|X_{t}}(\cdot|x_{0})$ in the right tail. Specifically, this can be established for kernel estimation with $r_{T}=b^{\text{dim}(X)/2}_{T}$, where $b_{T}$ is the bandwidth. The constant $c(x_{0})$ then becomes $\int K(u)^{2}du/g(x_{0})$ where $g(\cdot)$ denotes the density of $X$ and $K(\cdot)$ the kernel function (e.g., Cai, 2002; Li and Racine, 2008). [^3] These papers establish the pointwise convergence of the kernel conditional CDF estimator. Extension to the uniform convergence follows from the standard empirical process argument (e.g., Arcones and Yu, 1994). Such a result is available upon request. Assumption T describes the convergence of the tail index regression estimator. This condition is established by Nicolau et al. (2023) and Wang and Tsai (2009) for the case without subtracting $\bar{Y}$. Suppose $\bar{Y}$ converges to the median of $Y$ at the rate root-$T$, subtracting $\bar{Y}$ from the extreme values of $Y$ makes no asymptotic effect.
@@
Suppose Assumptions P, N, T, and R hold. Then,
@@ 5.1 Data Generating Designs
We develop two data generating designs based on the real data used by Adrian et al. (2019). These two designs are tailored to quarter-ahead predictions (Section 5.1.1) and year-ahead predictions (Section 5.1.2) of GDP growth rates, based on the current GDP growth rate and a financial indicator.
@@ 5.1.1 Design Based on Quarter-Ahead Predictions
These parameter values are obtained by fitting the known model of the conditional skewed-t distribution with the real data used by Adrian et al. (2019) for quarter-ahead predictions.
@@ 5.1.2 Design Based on Year-Ahead Predictions
These parameter values are obtained by fitting the known model of the conditional skewed-t distribution with the real data used by Adrian et al. (2019) for year-ahead predictions.
@@ 5.4 Additional Simulations
Having demonstrated the superior performance of our proposed method in predicting extreme growth events, we now apply this novel approach to a long-term U.S. macro-financial dataset spanning 1893–2016. The dataset is obtained from Gächter_Hasler_Huber_2025 stretching from 1893Q1 to 2016Q4 for the US and including annualized real GDP growth, a financial stress indicator, the 3-year average growth rate of the credit-to-GDP ratio, and the 3-year average growth rate of real house prices. The financial stress indicators is a historical newspaper-based financial stress indicator.
@@ 7 Summary and Discussions
While our primary focus is on the GaR framework, the proposed method is broadly applicable to other contexts involving the study of extreme outcomes, including Financial Stability at Risk (Adrian and Brunnermeier, 2016), Employment at Risk (Ramey and Zubairy, 2018), Growth at Risk (Adrian et al., 2019), Inflation at Risk (Loria et al., 2022), and Term Structure of Growth at Risk (Adrian et al., 2022), among others. We hope this paper will inspire further empirical research on these and other topics in economics and finance.
@@ Appendix A Mathematical Proofs
This DoA condition holds for a broad class of nonparametric distributions, including the skewed-t in particular. Thus, Adrian et al. (2019) implicitly assume this DoA condition too.
@@
This RV condition holds for a broad class of nonparametric distributions, including the skewed-t in particular. Thus, Adrian et al. (2019) implicitly assume this RV condition too.
@@ A.1 Proof of Proposition 1
Proof. For simplicity of writing, we abbreviate the notations $Y_{t+h}$ and $X_{t}$ as $Y$ and $X$, respectively, in this proof. We focus on the right tail, as the argument for the left tail follows symmetrically. Our proof strategy is outlined as follows: on the one hand, the skewed-t distribution implies the ratio of two conditional quantiles varies with $x$ through the exponent $v(x)$; on the other hand, a linear quantile regression structure implies that such a ratio remains constant over $x$, leading to the conclusion that $v(x)=\bar{v}$ for some constant $\bar{v}$.
We first investigate the DoA assumption. Specifically, since the skewed-t distribution satisfies the DoA assumption with $\xi\left(x\right)=1/v\left(x\right)>0$, it follows that $Q_{Y|X}\left(\tau|x\right)\sim c_{1}(x)(1-\tau)^{-1/v(x)}\rightarrow\infty$ as $\tau\rightarrow 1$. Therefore, we obtain that
$\frac{Q_{Y|X}\left(1-\frac{1}{t\tau}|x\right)}{Q_{Y|X}\left(1-\frac{1}{t}|x\right)}\sim\tau^{1/v(x)}$ (6)
for any $\tau>0$ as $t\to\infty$. Such a ratio changes over $x$ if and only if $v(x)$ is non-constant.
Now we explore the linear quantile regression structure. Let $x=(1,\dot{x}^{\prime})^{\prime}$ so that $Q_{Y|X}\left(\tau|x\right)=\alpha(\tau)+\dot{x}^{\prime}\beta(\tau)$. Let $j^{*}$ denote the coordinate of $\beta$ that has the heaviest tail, i.e., $\lim_{\tau\to 1}\beta_{j}(\tau)/\beta_{j^{*}}(\tau)=0$ for all $j\neq j^{*}$. [^4] If two components have the same order, i.e., $\lim_{\tau\to 1}\beta_{j}(\tau)/\beta_{j^{*}}(\tau)=c$ for some constant $c\in(0,\infty)$ and some $j$, then combine components $j$ and $j^{*}$ and consider $x_{j^{*}}+cx_{j}$ to be the dominating component. Now let $h=\lim_{\tau\to 1}\beta_{j^{*}}(\tau)/\alpha(\tau)\geq 0$. We now proceed with each of the three cases, $h=0$, $h=\infty$, and $h\in(0,\infty)$, separately.
Case 1: $h=0$. In this case, $\beta_{j^{*}}(1-1/t)=o(\alpha(1-1/t))$ as $t\rightarrow\infty$ and hence
@@
Case 3: $h\in(0,\infty)$. In this case, $h\alpha(1-1/t)\sim\beta_{j^{*}}(1-1/t)$ as $t\rightarrow\infty$ and hence
$\lim_{t\to\infty}\frac{Q_{Y|X}\left(1-\frac{1}{t\tau}|x\right)}{Q_{Y|X}\left(1-\frac{1}{t}|x\right)}=\lim_{t\to\infty}\frac{(\dot{x}_{j^{*}}+h^{-1})\beta_{j^{*}}\left(1-\frac{1}{t\tau}\right)}{(\dot{x}_{j^{*}}+h^{-1})\beta_{j^{*}}\left(1-\frac{1}{t}\right)}=\lim_{t\to\infty}\frac{\beta_{j^{*}}\left(1-\frac{1}{t\tau}\right)}{\beta_{j^{*}}\left(1-\frac{1}{t}\right)},$
@@
Combining three cases yields that $\lim_{t\to\infty}\frac{Q_{Y|X}\left(1-\frac{1}{t\tau}|x\right)}{Q_{Y|X}\left(1-\frac{1}{t}|x\right)}$ does not depend on $\dot{x}$. However, recall (6) that
$\frac{Q_{Y|X}\left(1-\frac{1}{t\tau}|x\right)}{Q_{Y|X}\left(1-\frac{1}{t}|x\right)}\rightarrow\tau^{1/v\left(x\right)}$